OpenStudy (anonymous):

what is 0.03 written in percent?

OpenStudy (anonymous):

3%

OpenStudy (anonymous):

3%

OpenStudy (anonymous):

3%

OpenStudy (anonymous):

You multiply 100 (%) by 0.03 :)

OpenStudy (anonymous):

0.03*100=3/100*100=3%

OpenStudy (anonymous):

ok i got it....

OpenStudy (anonymous):

just move the point 2 places right

OpenStudy (anonymous):

I think the answer has been aid many times :D